A woman has part of $4500 invested at 4% and the rest at 6%. If her annual return on each investment is the same, then what is the average rate of interest, which she realizes on the $4500?
Answer by atul(1)

Let A be invested at 4% and (4500 - A) at 6%
Interest at 4% is 0.04 * A
Interest at 6% is 0.06 (4500 - A)
Both are equal
0.04A = 0.06 * (4500 - A)
0.04A = 270 - 0.06 * A
0.1A = 270
A = 2700 and other investment is 4500-2700 = 1800